OpenStudy (anonymous):

A population of 1200 leopards decreases by 12% per year. How many leopards will there be in the population after 6 years? Round to the nearest whole number.

OpenStudy (anonymous):

is it a continouus decrease ? if it is in nature, most probably ?

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

Here P= 1200, r = 12%, n=6 Since population of tigers is depreciating, thus we use the formula: \[A= P(1-\frac{r}{100})^2\] Substituting the values: \[A= 1200(1-\frac{12}{100})^6\] \[= 1200(1-0.12)^6= 1200 (0.88)^6 =1200 \times 0.4644 =557.28 \approx 557\] Thus after 6 years approximately 557 tigers would be left. @blah124

OpenStudy (anonymous):

Thank you! Also for explaining it too.